OpenStudy (anonymous):

HELP!!!!!!!!!!! For the simple harmonic motion equation d=5sin(pi/3t) what is the period?

OpenStudy (anonymous):

period of \(\sin(bt) \) is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

\[\frac{2\pi}{\frac{\pi}{3}}=2\pi\times \frac{3}{\pi}=6\]